retired has been increasing over the last decade. This is?
The percentage of Americans who are retired has been increasing over the last decade. This is causing some concern because health care, social security, and other costs will be the responsibility of a smaller group of people. That is, as the percentage of retired people increases, the percentage of working age people decreases. The following model predicts the percentage of retired people based on demographic data:
R = t/873.36 - 2.15
where R is the percentage (as a decimal) of Americans who are retired in the year t. Use this model to complete the table below.
Year % of Retired People
blank 10%
blank 15%
blank 20%
I tried to rearrange the formula as: t=(R+2.15)873.36
but I'm still not getting the correct answers.

Answer by greenestamps(13195)



They should make it clear in the statement of the problem that the percentage R is a decimal.

If you are using 10, 15, and 20 as your R values, then you will get nonsensical answers. Use .10, .15, and .20.

(Your rearranged expression for t as a function of R is correct.)

-----------------------------------------------------------------------

(R+2.15)*873.36

10%: .10+2.15 = 2.25; 2.25*873.36 = 1965 to the nearest year
15%: .15+2.15 = 2.30; 2.30*873.36 = 2009
20%: .20+2.15 = 2.35; 2.35*873.36 = 2052

Are you not getting those numbers?